• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 15:57:17 on March 6,2009

    Entry number 264937

    keyword=Day Shift Summary

    Shift workers:

    Ole Hansen (SL)
    John LeRose (TO)
    Nuruzzaman (3rd)

    Summary:

    Finished 5-pass production with polarized 3He for kinematics G
    (p=1.20). Switched from 5 to 4-pass. Switched to longitudinal target
    polarization. Did beam checkout (spot/raster checks) and N2 reference
    cell runs for kinmematics 4I (p=1.42). Began polarized 3He production
    runs for kine 4I.

    Runs:

    LHRS:
    20561-20564 3He kin G production
    20565,20567 spot checks
    20566 junk, CODA problem
    20569 120 psig N2 ref cell kin 4I production
    20570 3He kin 4I production, in progress at end of shift

    BigBite:
    2013-2021 3He kin G
    2022 junk, MWDC HV not on
    2023 N2 kin 4I
    2024 3He kin 4I, in progress at end of shift